Exact travelling wave solutions for a class of nonlinear partial differential equations

Edamana Krishnan Sultan Qaboos University, Oman

The extended mapping method is developed to obtain the exact solutions to a nonlinear ordinary differential equation. As a simple application, a class of nonlinear partial differential equations such as nonlinear Klein-Gordon equation, a nonlinear evolution equation u_{tt} + αu_{xx} + βu + γu^{3} = 0, and a new coupled equation u_{xt} − 2uv = 0, v_{t} + 2uu_{x} = 0 are studied and multiple travelling wave solutions are obtained. Many new results are presented.